OpenStudy (anonymous):

indicate in standard form the equation of the line passing through the given point and having the given slope d(5,2) m=(2/5)

OpenStudy (texaschic101):

for this, we need point slope formula.. y - y1 = m(x - x1) (5,2) x1 = 5 and y1 = 2 slope(m) = 2/5 now we sub y - 2 = 2/5(x - 5) -- distribute through the parenthesis y - 2 = 2/5x - 2 -- add 2 to both sides y = 2/5x - 2 + 2 y = 2/5x + 0 standard form : Ax + By = C y = 2/5x + 0 -- no fractions, so multiply by 5 5y = 2x + 0 -- subtract 2x from both sides -2x + 5y = 0 -- multiply by -1 to make x positive 2x - 5y = 0 ==> standard form